Random variable X arises from a binomial experiment. Use n=8 and p=60 percent when answering the following:


a. Compute for the mean

b. Compute for the variance

c. Compute for the standard deviation


1
Expert's answer
2021-05-07T09:45:46-0400

a.

"mean=E(X)=np=8(0.6)=4.8"

b.


"Var(X)=\sigma^2=np(1-p)=8(0.6)(1-0.6)="

"=1.92"

c.


"\sigma=\sqrt{\sigma^2}=\sqrt{1.92}=0.8\sqrt{3}\approx1.3856"